how long is a spool of copper has a mass of 4.26 KG,diameter of 2.2 mm and a density of 8.96 g/cm3
1
Expert's answer
2013-03-12T12:24:21-0400
It is possible to find solution by using next formula:

m' = pi*r^2/(4*h'*d)

Where h' is 1 cm of length

Weight of 1 cm of length = pi (.22)^2/4 *1 *8.96 = 0.341 g

length of wire l = m/m' = 4260 g/0.341 = 12492cm or 124.9 m